What is the perimeter of a square which has the same area as a circle with circumference of 4π?
Found 2 solutions by stanbon, josgarithmetic:
Answer by stanbon(75887)

What is the perimeter of a square which has the same area as a circle with circumference of 4π?
-----
Perimeter of circle = pi*d = 4pi
Diameter = 4
Therefore radius = 2
--------------
Area of circle:: pi*r^2
Area = pi*2^2 = 4pi
----------------------------
Area of square = s^2 where s is the length of each side.
Equation:
s^2 = 4pi
s = 2*sqrt(pi)
-----------
Ans: Perimeter of square = 4*s = 4*2*sqrt(pi) = 8*sqrt(pi)
-------------------
Cheers,
Stan H.
----------

Answer by josgarithmetic(39617)

2pi*r=4pi
r=(4pi)/(2pi)
r=2
-
s=pi*r^2
s=pi*2^2
s=4pi
-


x, x, x, x, total sides of the square.
x^2=s
x^2=4pi
x=2*sqrt(pi)

-
-
4*2*sqrt(pi)


ANSWER:
-------------------
8*sqrt(pi)
-------------------